OpenStudy (m_m13):

one question will fan and medal What is the solution to the equation? 2x^2+4=20 choose all the answers that are correct.

OpenStudy (m_m13):

A.

OpenStudy (m_m13):

B.

OpenStudy (m_m13):

C.

OpenStudy (m_m13):

D.

OpenStudy (m_m13):

@ILovePuppiesLol

OpenStudy (anonymous):

None, lol

OpenStudy (anonymous):

\[2\sqrt{2} \] would work tho

OpenStudy (anonymous):

subtract 4 from each side 2x^2 = 16 divide each side by 2 x^2 = 8 x= sqrt8